# Count on Your Elders: Laplace vs Gaussian Noise

### v2

@@ Abstract
First, we consider the problem of counting under continual observation and present a new generalization of the binary tree mechanism that uses a $k$-ary number system with negative digits to improve the privacy-accuracy trade-off. Our mechanism uses Laplace noise and improves the mean squared error over all “optimal” $(\varepsilon,\delta)$-differentially private factorization mechanisms based on Gaussian noise whenever $\delta$ is sufficiently small. Specifically, using $k=19$ we get an asymptotic improvement over the bound given in the work by Henzinger, Upadhyay and Upadhyay (SODA 2023) when $\delta=O(T^{-0.92})$.
Second, we show that the noise added by the Gaussian mechanism can always be replaced by Laplace noise of comparable variance for the same $(\varepsilon,\delta)$ privacy guarantee, and in fact for sufficiently small $\delta$ the variance of the Laplace noise becomes strictly better. This challenges the conventional wisdom that Gaussian noise should be used for high-dimensional noise.
@@ 1 Introduction
While continual counting is a deceptively simple problem, we note that it still has signifcant gaps between upper and lower bounds [16, 24, 19, 17, 9]. The problem is used as a subroutine in various applications ranging from private learning [32, 28, 8, 12, 7] to histogram estimation [4, 5, 27, 36], so any improvement in upper bounds affect these applications.
@@ 1.1 Our contributions
Here and elsewhere $\log_{k}$ is the base $k$ logarithm, and $\log(\cdot)=\log_{2}(\cdot)$. To the best of our knowledge, this is the best bound on the mean squared error known for any $\varepsilon$-DP mechanism, up to lower-order asymptotically vanishing terms. It improves the error achieved by applying techniques in [26] to the binary tree mechanism (as implemented in [28] and referred to elsewhere as ?Honaker Online?) by a factor 4 when $k=19$. While this is interesting in its own right, we note that this improvement in the leading constant has implications for how the error compares to $(\varepsilon,\delta)$-DP mechanisms. Henzinger, Upadhyay and Upadhyay [24] produce a factorization with a bound on the mean squared error of $C_{\varepsilon,\delta}^{2}\big{(}1+\frac{\ln(4T/5)}{\pi}\big{)}^{2}$, where $C_{\varepsilon,\delta}=\frac{2}{\varepsilon}\sqrt{\frac{4}{9}+\ln\big{(}\frac{1}{\delta}\sqrt{\frac{2}{\pi}}\big{)}}$, and show that no factorization mechanism based on Gaussian noise can asymptotically improve on this error. [^1] As the constant $C_{\varepsilon,\delta}$ is valid for all $\varepsilon,\delta\in(0,1)$, it seems plausible that it can be reduced via a tighter analysis of the Gaussian mechanism, as done in for example [2]. We improve on their error using Laplacian noise for small enough $\delta$.
@@
Given an integer $T\geq 2$, there exists an $\varepsilon$-DP algorithm for continual counting, based on Laplacian noise, that improves on the mean squared error bound given in [24], up to lower-order asymptotically vanishing terms, whenever $\delta=O(T^{-0.92})$.
@@ 1.2 Overview of technical ideas
We start out by building a complete $k$-ary tree of height $h=\lceil\log_{k}(T+1)\rceil$ on top of the input $\mathbf{x}\in\{0,1\}^{T}$, see Figure 1. To analyze this, it turns out that expressing a time step $t$ as a $k$-ary number, i.e., as $h$ digits in $\{0,\dots,k-1\}$, i.e. $\mathbf{t}\in\{0,\dots,k-1\}^{h}$, gives a direct means of computing which vertices contribute to a given prefix sum. Increasing the arity of the tree reduces the height, which reduces the $\ell_{1}$-sensitivity and levels of vertices that might get added, but increases the number of vertices to add per level. To mitigate the increase in number of vertices to add, we leverage subtraction, see Figure 2
@@
To derive the $(\varepsilon,\delta)$-DP guarantee of the Laplace mechanism, we make use of a heterogeneous composition theorem [29]. Given two arbitrary multidimensional neighboring inputs, we argue for the privacy loss in each coordinate when one of them is released with respect to its neighbor. To get the aggregate privacy guarantee for the entire vector, we apply the composition theorem over the individual coordinates being released, and show that the guarantee must hold for any two neighboring outputs. It turns out that, as in the case of the Gaussian mechanism, the $(\varepsilon,\delta)$-DP guarantee for Laplacian noise naturally depends on the $\ell_{2}$-sensitivity of the function being released.
@@ 1.3 Related work
Historically, the first mechanisms with good error are typicallywcredited to Chan, Shi, and Song [6] and Dwork, Naor, Pitassi, and Rothblum [16], though similar constructions were proposed independently by others [23, 20]. These results are commonly referred to as binary tree mechanisms. Honaker [26] observed that the full tree used by the binary tree mechanism can be used to produce more efficient estimates since each prefix query can be computed using different combinations of vertices. A subset of Honaker’s techniques can be used to reduce the mean squared error by up to a factor 2, at some cost in efficiency. Rather than leveraging the density of the tree to improve the error, Andersson and Pagh [1] identified that making the tree sparser – only using a subset of the leaves to store inputs – allowed for a factor 2 reduction in the error while keeping the variance of outputs constant, for $(\varepsilon,\delta)$-DP. Although they do not state a result for it, using their technique for $\varepsilon$-DP yields a factor 4 improvement over the binary tree mechanism, which is still more than a factor 2 away from our improvement. We also mention the work of Qardaji, Yang, and Li [33] which, while focusing on range queries, did investigate reducing the mean squared error by means of considering trees of higher arity. They claim an optimal arity of $k=16$, but average over a different set of queries, and do not leverage the subtraction of vertices. Therefore, we emphasize that $k$-ary trees for range queries have already been studied. However, to the best of our knowledge, the connection between $k$-ary numbers and the variance of outputs have not been explicitly made, nor does there seem to be any work on $k$-ary trees for prefix sums in particular.
Many recent papers [14, 12, 24, 19] have focused on improving the error for continual counting for $(\varepsilon,\delta)$-DP by directly searching among factorizations of $A$, the lower-triangular all-1s matrix. Most notably, Henzinger, Upadhyay, and Upadhyay [24] showed that there exists an optimal factorization $L=R=\sqrt{A}$ (also identified by Bennett [3]) whose error is optimal up to the leading constant across all factorizations. But, importantly, this is for $(\varepsilon,\delta)$-DP and assumes using Gaussian noise. Due to the high $\ell_{1}$-sensitivity of $R=\sqrt{A}$, the noise needed to release $R\mathbf{x}$ with the (conventional) Laplacian mechanism results in an $\varepsilon$-DP matrix mechanism with worse error than that of the binary mechanism.
@@ 2 Preliminaries
As the main goal of this paper is the study and design of differentially private algorithms for counting queries, we will first formalize the problem. Let vector $\mathbf{x}\in\{0,1\}^{T}$ be an input stream, where $\mathbf{x}_{t}$ is received at time $t$. Then, two input streams $\mathbf{x},\mathbf{x}^{\prime}$ are neighbouring inputs, denoted $\mathbf{x}\sim\mathbf{x}^{\prime}$, if they take on the same value at every time step except one. The target function we wish to privately release is $f:\{0,1\}^{T}\to\mathbb{R}$ where $f(\mathbf{x})_{t}=\sum_{i=1}^{t}\mathbf{x}_{i}$. We say that a random algorithm :$\mathcal{M}:\{0,1\}^{T}\to\mathbb{R}^{T}$ is a differentially private mechanism for continual counting if it satisfies either $\varepsilon$-DP or $(\varepsilon,\delta)$-DP under this neighbouring relation.
@@ 3.1 Binary Tree Mechanisms
The first algorithms [6, 16, 20, 23] with good utility are based on binary trees and are referred to as the binary tree mechanism. Implementation details vary, but we will consider a formulation of it where only the left subtrees are used, and in particular not the root vertex, structurally most similar to Algorithm 2 in [6]. The main idea for the algorithm is to build a binary tree of height $h=\lceil\log(T+1)\rceil$ on top of the input stream $\mathbf{x}$ where each leaf stores an input $\mathbf{x}_{t}$, and each vertex higher up in the tree stores the sum of its children. Having such a tree, it is possible to add together vertices in the tree in such a way that at most $1$ vertex per level in the tree is used, up to $h$ in total, where the sum corresponds to a prefix sum up to $t\in[1,T]$. In particular, which vertices that get summed to produce a given prefix sum is described exactly by the $h$-bit binary representation of the current time step, where each $1$ corresponds to adding a vertex in the tree to the output. To make this private, we add i.i.d. Laplacian noise $\mathsf{Lap}(\Delta_{1}/\varepsilon)$ to each vertex in the tree, where $\Delta_{1}=h$ since the trees for two neighbouring inputs will differ in one vertex per level, excluding the root which is not used in computations. We defer making exact statements about the corresponding utility to Section 3.2, where this description of the binary tree mechanisms is a special case of Algorithm 1 for $k=2$.
@@ 3.2 $k$-ary Tree Mechanisms
We will next consider trees of greater arity $k\geq 2$ to improve constants for the error. We note that using trees of higher arity has already been studied [33, 10], but there applied to the analysis of range queries, of which prefix sums are a special case. With a greater arity tree, a lesser height is needed to accommodate a given number of leaves. However, it might also require adding more vertices per level to generate a given prefix sum estimate. To explore this trade-off, we introduce notation for $k$-ary trees next.
@@
Note that line 11 always adds the noisy count from a vertex in the tree at level $\ell\in[1,h]$ since $p=p_{n}\overset{\pmod{k^{\ell-1}}}{=}0$, and since $\mathbf{t}$ is a valid representation of a time step $t\leq T$, we also have $p\leq t$, and so the vertex must be contained in the tree. Note further that everything after line 4 can be considered post-processing of privately releasing the tree $\mathcal{T}_{k,h}$ storing partial sums. Since we never use the root vertex of the tree at level $h+1$, any two neighbouring inputs will differ at exactly $h$ vertices which we use, therefore $\Delta_{1}=h$, implying that releasing the individual counts in the tree with $\mathsf{Lap}(h/\varepsilon)$ gives privacy.
@@
Identifying that the mean squared error over $T$ steps is equal to the average number of vertices added together to produce an output, call this number $a_{[1,T]}$, times the variance of each vertex, $2h^{2}/\varepsilon^{2}$. Observe that $a_{[1,T]}=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname{\mathsf{rep}}_{k}(t,h)\right\rVert_{1}$ and that $T$ is the greatest integer possible to represent with $h$ digits from $\{0,1,\dots,k-1\}$. This implies that we can compute the average $\ell_{1}$-weight over $\mathbf{r}\in\{0,1,\dots,k-1\}^{h}$, $a_{[0,T]}=h\cdot\frac{1}{k}\sum_{d=0}^{k-1}|d|=\frac{h(k-1)}{2}$. Since $T\cdot a_{[1,T]}=(T+1)\cdot a_{[0,T]}$, we have that $a_{[1,T]}=(1+1/T)\cdot a_{[0,T]}=\frac{h(k-1)}{2(1-1/k^{h})}$. Plugging this in gives a mean squared error of $\frac{(k-1)h}{2(1-1/k^{h})}\cdot\frac{2h^{2}}{\varepsilon^{2}}=\frac{(k-1)h^{3}}{\varepsilon^{2}(1-1/k^{h})}$.
@@
Given integer $T\geq 2$ and constant integer $k\geq 2$, there exists an $\varepsilon$-DP mechanism for continual counting that on receiving a stream $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ achieves a mean squared error of $\frac{(k-1)h}{\varepsilon^{2}\log(k)^{3}}\cdot\log(T)^{3}+o(\log(T)^{3})$, computes all outputs in time $O(T)$ and $O(\log T)$ space.

### v3

@@ Abstract
First, we consider the problem of counting under continual observation and present a new generalization of the binary tree mechanism that uses a $k$-ary number system with negative digits to improve the privacy-accuracy trade-off. Our mechanism uses Laplace noise and whenever $\delta$ is sufficiently small it improves the mean squared error over the best possible $(\varepsilon,\delta)$-differentially private factorization mechanisms based on Gaussian noise. Specifically, using $k=19$ we get an asymptotic improvement over the bound given in the work by Henzinger, Upadhyay and Upadhyay (SODA 2023) when $\delta=O(T^{-0.92})$.
Second, we show that the noise added by the Gaussian mechanism can always be replaced by Laplace noise of comparable variance for the same $(\varepsilon,\delta)$-differential privacy guarantee, and in fact for sufficiently small $\delta$ the variance of the Laplace noise becomes strictly better. This challenges the conventional wisdom that Gaussian noise should be used for high-dimensional noise.
Finally, we study whether counting under continual observation may be easier in an average-case sense than in a worst-case sense. We show that, under pure differential privacy, the expected worst-case error for a random input must be $\Omega(\log(T)/\varepsilon)$, matching the known lower bound for worst-case inputs.
@@ 1 Introduction
While continual counting is a deceptively simple problem, we note that it still has significant gaps between upper and lower bounds [16, 25, 19, 17, 9]. The problem is used as a subroutine in various applications ranging from private learning [33, 29, 8, 12, 7] to histogram estimation [4, 5, 28, 37], so any improvement in upper bounds affect these applications.
@@ 1.1 Our contributions
We denote the base $k$ logarithm by $\log_{k}$, and define $\log(\cdot)=\log_{2}(\cdot)$. To the best of our knowledge, this is the best bound on the mean squared error known for any $\varepsilon$-DP mechanism, up to lower-order asymptotically vanishing terms. It improves the error achieved by applying techniques in [27] to the binary tree mechanism (as implemented in [29] and referred to elsewhere as ?Honaker Online?) by a factor 4 when $k=19$. While this is interesting in its own right, we note that this improvement in the leading constant has implications for how the error compares to $(\varepsilon,\delta)$-DP mechanisms. Henzinger, Upadhyay and Upadhyay [25] produce a factorization with a bound on the mean squared error of $C_{\varepsilon,\delta}^{2}\big{(}1+\frac{\ln(4T/5)}{\pi}\big{)}^{2}$, where $C_{\varepsilon,\delta}=\frac{2}{\varepsilon}\sqrt{\frac{4}{9}+\ln\big{(}\frac{1}{\delta}\sqrt{\frac{2}{\pi}}\big{)}}$, and show that no factorization mechanism based on Gaussian noise can asymptotically improve on this error. [^1] As the constant $C_{\varepsilon,\delta}$ is valid for all $\varepsilon,\delta\in(0,1)$, it seems plausible that it can be reduced via a tighter analysis of the Gaussian mechanism, as done in for example [2]. We improve on their error using Laplace noise for small enough $\delta$.
@@
Given an integer $T\geq 2$, there exists an $\varepsilon$-DP algorithm for continual counting, based on Laplace noise, that improves on the mean squared error bound given in [25], up to lower-order asymptotically vanishing terms, whenever $\delta=O(T^{-0.92})$.
@@
Finally, in Section 6 we show the following lower bound, which matches the classical packing lower bound for continual observation under pure differential privacy [16, 18] in the case where the input is random.
Theorem 1.4.
Let $\mathcal{U}$ be the distribution of strings of length $T$ where every element in the string is drawn from $B(1/2)$ (Bernoulli distribution with $p=1/2$). Let $\mathcal{M}$ be an $\varepsilon$-differentially private algorithm that solves the continual counting problem for strings from $\mathcal{U}$ with error at most $\alpha$ with probability at least $3/4$ (probability over random input and random choices of the algorithm). Then $\alpha=\Omega(\varepsilon^{-1}\log T)$.
@@ 1.2 Overview of technical ideas
Improved tree-based mechanism
We start out by building a complete $k$-ary tree of height $h=\lceil\log_{k}(T+1)\rceil$ on top of the input $\mathbf{x}\in\{0,1\}^{T}$, see Figure 1. To analyze this, it turns out that expressing a time step $t$ as a $k$-ary number, i.e., as $h$ digits in $\{0,\dots,k-1\}$, i.e. $\mathbf{t}\in\{0,\dots,k-1\}^{h}$, gives a direct means of computing which vertices contribute to a given prefix sum. Increasing the arity of the tree reduces the height, which reduces the $\ell_{1}$-sensitivity and levels of vertices that might get added, but increases the number of vertices to add per level. To mitigate the increase in number of vertices to add, we leverage subtraction, see Figure 2.
@@
Factorization mechanism with Laplace noise
To derive the $(\varepsilon,\delta)$-DP guarantee of the Laplace mechanism (Theorem 1.3), we make use of a heterogeneous composition theorem [30]. Given two arbitrary multidimensional neighboring inputs, we argue for the privacy loss in each coordinate when one of them is released with respect to its neighbor. To get the aggregate privacy guarantee for the entire vector, we apply the composition theorem over the individual coordinates being released, and show that the guarantee must hold for any two neighboring outputs. It turns out that, as in the case of the Gaussian mechanism, the $(\varepsilon,\delta)$-DP guarantee for Laplace noise naturally depends on the $\ell_{2}$-sensitivity of the function being released.
Lower bound for random inputs
Finally, the lower bound of Theorem 1.4 is shown by a variation of the packing argument that has been used to show an $\Omega(\log(T)/\varepsilon)$ lower bound for worst-case inputs. The core idea is to construct a set of correlated input strings that are individually close to uniformly random, but a continual observation mechanism that has low maximum error on all of them would be able to reliably distinguish these input strings. More precisely, for integers $T$, $B$, $m=T/B$ and $k\leq B/4$, we construct our collection of strings $\mathbf{x}^{(0)},\mathbf{x}^{(1)},\dots,\mathbf{x}^{(m)}$ by first decomposing $[T]$ into blocks of size $B:B_{1}=[1,B],B_{2}=[B+1,2B],\dots,B_{m}=[T-B+1,T]$ and then doing the following sampling process: $\mathbf{x}^{(0)}$ is sampled uniformly from $\{0,1\}^{T}$ conditioned on each block $B_{i}$ containing between $B/4$ and $43B/$ 4 1s, and for $i\in[m]$, $\mathbf{x}^{(i)}$ is derived from $\mathbf{x}^{(0)}$ by flipping $k$ 0s to 1s in block $B_{i}$. We show that each individual string has $o(1)$ total variation distance to the uniform distribution over $\{0,1\}^{T}$, and thus a mechanism accurate on the uniform distribution will also be accurate on our random strings. Also, pairs of our random strings will be $2k$-neighboring with probability 1. The key insight is that the difference of the true count of $\mathbf{x}^{(i)}$ minus $\mathbf{x}^{(0)}$ is $k$ at the end of $B_{i}$, and zero at the end of each preceding $B_{j}$ for $j<i$, and this event is detectable if the mechanism is sufficiently accurate on both inputs. Choosing parameters carefully, a packing-like argument in the end shows that a too low error on uniformly distributed inputs contradicts differential privacy.
@@ 1.3 Related work
Historically, the first mechanisms with good error are typically credited to Chan, Shi, and Song [6] and Dwork, Naor, Pitassi, and Rothblum [16], though similar constructions were proposed independently by others [23, 20]. These results are commonly referred to as binary tree mechanisms. Honaker [27] observed that the full tree used by the binary tree mechanism can be used to produce more efficient estimates since each prefix query can be computed using different combinations of vertices. A subset of Honaker’s techniques can be used to reduce the mean squared error by up to a factor 2, at some cost in efficiency. Rather than leveraging the density of the tree to improve the error, Andersson and Pagh [1] identified that making the tree sparser – only using a subset of the leaves to store inputs – allowed for a factor 2 reduction in the error while keeping the variance of outputs constant, for $(\varepsilon,\delta)$-DP. Although they do not state a result for it, using their technique for $\varepsilon$-DP yields a factor 4 improvement over the binary tree mechanism, which is still more than a factor 2 away from our improvement. We also mention the work of Qardaji, Yang, and Li [34] which, while focusing on range queries, did investigate reducing the mean squared error by means of considering trees of higher arity. They claim an optimal arity of $k=16$, but average over a different set of queries, and do not leverage the subtraction of vertices. Cardoso and Rogers [4] used $k$-ary trees for prefix sums as a subroutine for $(\varepsilon,\delta)$-DP and investigated choosing $k$ to minimize the maximum variance.
Many recent papers [14, 12, 25, 19] have focused on improving the error for continual counting for $(\varepsilon,\delta)$-DP by directly searching among factorizations of $A$, the lower-triangular all-1s matrix. Most notably, Henzinger, Upadhyay, and Upadhyay [25] showed that there exists an optimal factorization $L=R=\sqrt{A}$ (also identified by Bennett [3]) whose error is optimal up to the leading constant across all factorizations. But, importantly, this is for $(\varepsilon,\delta)$-DP and assumes using Gaussian noise. Due to the high $\ell_{1}$-sensitivity of $R=\sqrt{A}$, the noise needed to release $R\mathbf{x}$ with the (conventional) Laplace mechanism results in an $\varepsilon$-DP matrix mechanism with worse error than that of the binary mechanism.
@@
Finally, we turn to lower bounds where the lower bound of [16] remains the best bound known on the noise needed for continual observation under pure differential privacy. The hard inputs used in the proof (see [18] for details) are very sparse vectors with a single block of 1s that can be reliably distinguished by any algorithm that outputs prefix sums with noise significantly smaller than the block size, yet are close to each other in neighbor distance. This structure means that a “packing argument” can be used to show a lower bound on the amount of noise needed. One might wonder if most inputs are easier and allow a smaller error — our lower bound shows that this is not the case. Some algorithmic problems have the property that worst-case problem instances can be reduced to random instances (for example [24]). Though we are not aware of such a reduction for continual counting, our lower bound can be seen as an instantiation of this general principle.
@@ 2 Preliminaries
As the main goal of this paper is the study and design of differentially private algorithms for counting queries, we will first formalize the problem. Let vector $\mathbf{x}\in\{0,1\}^{T}$ be an input stream, where $\mathbf{x}_{t}$ is received at time $t$. Then, two input streams $\mathbf{x},\mathbf{x}^{\prime}$ are neighboring inputs, denoted $\mathbf{x}\sim\mathbf{x}^{\prime}$, if they take on the same value at every time step except one. The target function we wish to privately release is $f:\{0,1\}^{T}\to\mathbb{R}^{T}$ where $f(\mathbf{x})_{t}=\sum_{i=1}^{t}\mathbf{x}_{i}$. We say that a random algorithm :$\mathcal{M}:\{0,1\}^{T}\to\mathbb{R}^{T}$ is a differentially private mechanism for continual counting if it satisfies either $\varepsilon$-DP or $(\varepsilon,\delta)$-DP under this neighboring relation.
@@ 3.1 Binary Tree Mechanisms
The first algorithms [6, 16, 20, 23] with good utility are based on binary trees and are referred to as the binary tree mechanism. Implementation details vary, but we will consider a formulation of it where only the left subtrees are used, and in particular not the root vertex, structurally most similar to Algorithm 2 in [6]. The main idea for the algorithm is to build a binary tree of height $h=\lceil\log(T+1)\rceil$ on top of the input stream $\mathbf{x}$ where each leaf stores an input $\mathbf{x}_{t}$, and each vertex higher up in the tree stores the sum of its children. Having such a tree, it is possible to add together vertices in the tree in such a way that at most $1$ vertex per level in the tree is used, up to $h$ in total, where the sum corresponds to a prefix sum up to $t\in[1,T]$. In particular, which vertices that get summed to produce a given prefix sum is described exactly by the $h$-bit binary representation of the current time step, where each $1$ corresponds to adding a vertex in the tree to the output. To make this private, we add i.i.d. Laplace noise $\mathsf{Lap}(\Delta_{1}/\varepsilon)$ to each vertex in the tree, where $\Delta_{1}=h$ since the trees for two neighboring inputs will differ in one vertex per level, excluding the root which is not used in computations. We defer making exact statements about the corresponding utility to Section 3.2, where this description of the binary tree mechanisms is a special case of Algorithm 1 for $k=2$.
@@ 3.2 $k$-ary Tree Mechanisms
We will next consider trees of greater arity $k\geq 2$ to improve constants for the error. We note that using trees of higher arity has already been studied [34, 10, 4]. With a greater arity tree, a lesser height is needed to accommodate a given number of leaves. However, it might also require adding more vertices per level to generate a given prefix sum estimate. To explore this trade-off, we introduce notation for $k$-ary trees next.
@@
Note that line 11 always adds the noisy count from a vertex in the tree at level $\ell\in[1,h]$ since $p=p_{n}\overset{\pmod{k^{\ell-1}}}{=}0$, and since $\mathbf{t}$ is a valid representation of a time step $t\leq T$, we also have $p\leq t$, and so the vertex must be contained in the tree. Note further that everything after line 4 can be considered post-processing of privately releasing the tree $\mathcal{T}_{k,h}$ storing partial sums. Since we never use the root vertex of the tree at level $h+1$, any two neighboring inputs will differ at exactly $h$ vertices which we use, therefore $\Delta_{1}=h$, implying that releasing the individual counts in the tree with $\mathsf{Lap}(h/\varepsilon)$ gives privacy.
@@
Identifying that the mean squared error over $T$ steps is equal to the average number of vertices added together to produce an output, call this number $a_{[1,T]}$, times the variance of each vertex, $2h^{2}/\varepsilon^{2}$. Observe that $a_{[1,T]}=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname{\mathsf{rep}}_{k}(t,h)\right\rVert_{1}$ and that $T$ is the greatest integer possible to represent with $h$ digits from $\{0,1,\dots,k-1\}$. This implies that we can compute the average $\ell_{1}$-weight over $\{0,1,\dots,k-1\}^{h}$, $a_{[0,T]}=h\cdot\frac{1}{k}\sum_{d=0}^{k-1}|d|=\frac{h(k-1)}{2}$. Since $T\cdot a_{[1,T]}=(T+1)\cdot a_{[0,T]}$, we have that $a_{[1,T]}=(1+1/T)\cdot a_{[0,T]}=\frac{h(k-1)}{2(1-1/k^{h})}$. Plugging this in gives a mean squared error of $\frac{(k-1)h}{2(1-1/k^{h})}\cdot\frac{2h^{2}}{\varepsilon^{2}}=\frac{(k-1)h^{3}}{\varepsilon^{2}(1-1/k^{h})}$.
@@
Given integer $T\geq 2$ and constant integer $k\geq 2$, there exists an $\varepsilon$-DP mechanism for continual counting that on receiving a stream $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ achieves a mean squared error of $\frac{k-1}{\varepsilon^{2}\log(k)^{3}}\cdot\log(T)^{3}+o(\log(T)^{3})$, computes all outputs in time $O(T)$ and $O(\log T)$ space.